OpenStudy (anonymous):

Katrina has invested some money for 12 years at a fixed interest rate that is compounded continuously. If her money will double in 8 years. will it triple in 12 years?

OpenStudy (rsadhvika):

given that money will double in 8 years \(2 = e^{8r} \\ => r = \frac{\ln 2}{8} ----- (1)\) need to check if it triples in 12 years \(3 = e^{12r} \\ => r = \frac{\ln 3}{12} ----- (2)\) combining (1) and (2) \(\frac{\ln 2}{8}= \frac{\ln 3}{12} \) \(\frac{\ln 2}{\ln 3}= \frac{8}{12} \) \(\frac{\ln 2}{\ln 3}= \frac{2}{3} \) evaluate the log, if this true then the money will triple in 12 years